OpenStudy (anonymous):

I need help proving trigonometric identities. Example: (cot x sin x)(sec x – cos x) = sin2 x

sam (.sam.):

\[\Large (=\frac{cosx}{sinx}sinx)(\frac{1}{cosx}-cosx) \\ \\ \Large =cosx(\frac{1}{cosx}-cosx) \\ \\ \Large =1-\cos^2x \] Then use an identity

OpenStudy (anonymous):

How do I find the reason for each calculation?

OpenStudy (goformit100):

For Solving this Question, I would like to ask you that, do you have the detailed conception of : Trigonometric Identity.

OpenStudy (anonymous):

@goformit100 No :(

sam (.sam.):

You need to simplify it, and as always, you want to change trigonometry such as cotx, secx, and cscx into sinx/cosx , 1/cosx, and 1/sinx

OpenStudy (unklerhaukus):

to start the problems just convert each trig function to sines and cosines, as .sam. has done in first step

OpenStudy (anonymous):

So change cot x to cos x/sin x and sec x to 1/cos x then why do you cancel sin x?

OpenStudy (anonymous):

Also do reasons need to meet a certain standard, I know there needs to be four of them in this case but I don't know why?

OpenStudy (anonymous):

@.Sam. Sorry to tag you but I thought you might know the answer to my response question.

sam (.sam.):

You have to cancel in order to simplify it

OpenStudy (anonymous):

And how is cos x((1 / cos x) – cos x) able to change to (cos x / cos x) – cos2 x?

sam (.sam.):

You distribute

sam (.sam.):